# Using matrix-product states for time-series machine learning

### v1

@@ Appendix D Imputation algorithm details
Here we provide details on the specific steps of the imputation algorithm used in the main text. To elucidate the key steps of the proposed algorithm, consider a 6 site MPS trained on time-series instances of length $T=6$ samples from class C, shown in Fig. 11. Given an unseen time-series instance of data class C that is incomplete, for example, only the values $x_{2}$, $x_{4}$, $x_{6}$ are known, the goal of imputation is to estimate the unknown values (i.e., to ‘fill in the gaps’) - in this case, $x_{1}$, $x_{3}$, $x_{5}$. Our approach is performed in two main steps: (i) the trained MPS is conditioned on known values by first transforming them into quantum states and then projecting the MPS onto these known states; (ii) using the conditioned MPS, the unknown values are then determined by sequentially computing and sampling from a series of single-site reduced density matrices.
Figure 11: A toy imputation problem involving six time-series samples. Here, we consider three observed data points (solid pink markers) $x_{2}$, $x_{4}$, $x_{6}$, and three unobserved (missing) data points, $x_{1}$, $x_{3}$, $x_{5}$. Observed values in the encoding domain $x_{t}\in[a,b]$ are mapped to discrete quantum states (pink vectors) via the encoding $\phi_{i}(x_{i})$ i.e., $s_{i}=\phi_{i}(x_{i})$. The MPS is then conditioned on the observed values by making a series of projective measurements at the sites corresponding to the observed states, $s_{2}$, $s_{4}$, $s_{6}$. The unobserved states (orange vectors) are determined from the conditional probability distribution encoded by the partially conditioned MPS, and transformed back to time-series values in the encoding domain (orange markers) by an inverse transformation.
@@ D.1 Projecting the trained MPS onto known states
The first step of the algorithm is to project the MPS - through a sequence of iterative site-wise measurements - into a subspace which can then be used to impute missing data. We assume the MPS $W$, which is trained to approximate the underlying joint PDF of class C time-series data, is L2 normalized. This normalization condition can be shown diagrammatically in Penrose graphical notation [49]) as:
@@
Here, we show the resulting 4 site MPS which approximates the PDF conditional on having measured the states $s_{2},s_{4}$.
@@ D.2 Imputing missing time-series data
The second step of the imputation algorithm involves determining the values of the unknown data points by estimating their respective values from the conditional PDF represented by the MPS. Here we use a sequential sampling approach which involves computing $d\times d$ single-site reduced density matrices.
@@
which ensures the pdf remains normalized over the support of the encoding domain $[a,b]$ ($[-1,1]$ for Legendre). For details about our choice of the median as the best single point estimate, see Appendix. B.
@@
To transform the imputed states (yellow vectors) back to their rescaled time-series values, one can apply the inverse of the feature map $\phi_{t}^{-1}(x)$. An additional transformation can be applied to convert the rescaled time-series (in the encoding domain) back to their respective amplitudes in the original data domain, i.e., .
@@ Appendix E Inverse transform sampling
To generate single-shot time-series trajectories, as in Fig. 7, we used an inverse transform sampling (ITS) approach augmented with a rejection scheme to mitigate the effects of the encoding error discussed in Sec. B. Specifically, due to the finite basis approximation of (original continuous) time-series values, we observe an error which manifests as an artificial broadening of the conditional distribution of time-series values at each MPS site. As a consequence of this broadening, the conditional distribution can assign non-negligible probabilities to time-series values that would otherwise be highly improbable, as depicted in Fig. 9. When generating trajectories through conditional site-wise sampling – starting at the first site and proceeding sequentially to the end – these errors can propagate and compound, potentially leading to trajectories that are not representative of the underlying joint distribution. While increasing $d$, the number of basis functions in the encoding (as in Eq. (3)), can mitigate this error in principle, achieving meaningful reductions in distribution broadening would require values that render ITS computationally intractable for most applications. Instead, we use a rejection scheme to discard highly improbable values based on their deviation from the median (WMAD) of the conditional distribution. Our choice of the median was motivated empirically from the analysis presented in Appendix B.
@@
Using the inverse cumulative distribution function, $F_{i}^{-1}(u)$, we select the value $x_{i}$ such that $F_{i}(x_{i})=u$. We then apply the WMAD-based rejection criteria described in Alg.1.
The selection of the rejection threshold factor $\alpha$ requires careful consideration. A threshold that is overly restrictive (i.e., small $\alpha$) would constrain the sampling space, potentially introducing bias by suppressing variability that is a feature of the true encoded distribution. Conversely, a less stringent threshold (i.e., large $\alpha$) would fail to adequately address the artificial broadening effects. Through empirical investigation, we identified $\alpha=2.0$ (i.e., $2\cdot\rm WMAD)$ as an effective heuristic, which provides a balance between the two extremes.
Algorithm 1 Single-shot time-series generation.
@@
Appendix F Real world time-series datasets
Here we provide further details on the three UCR time-series datasets used for the classification task in Sec.III.3, each representing key domain of application: (i) ECG (medicine), (ii) Power Demand (energy) and (iii) Astronomy (physics). Representative time-series examples from each dataset are shown in Fig. 12.
Figure 12: Representative time-series instances for each empirical dataset. Several time-series instances from each class are plotted on the same axis, and in the original data domain, for the three datasets: (a) Astronomy (2-class subset of KeplerLightCurves), (b) ECG (ECG200), and (c) Power Demand (ItalyPowerDemand). In each panel, we plot representative instances from (i) class 0, and (ii) class 1, for the various binary classification problems.
Figure 13: MPS imputations on unseen Astronomy time series with real data gaps caused by satellite transmission downlinks. The panels depict examples of MPS ($\chi_{\rm max}=35$, $d=12$) and 1-NNI on held-out instances with real data gaps from four classes of star system: (a) RR-Lyrae variable, (b) contact eclipsing binary (contact EB), (c) non variable, and (d) $\gamma$ Doradus variable. The flat horizontal lines in light gray correspond to real missing data which has been imputed with the time-series mean in the original UCR dataset.
@@ F.1 Imputation on real-world data gaps
For our investigation of the Astronomy dataset in Sec. III.2 of the main text, we excluded time-series windows with true data gaps for which ground-truth values are unavailable. For completeness, to showcase how our method may perform with real missing data, we trained a $d=12$, $\chi_{\rm max}=35$ MPS on clean windows from four classes of star system: (i) RR-Lyrae variable; (ii) contact Eclipsing Binary (contact EB); (iii) non variable; and, (iv) Gamma Doradus variable ($\gamma$ Doradus variable). We then imputed real missing values on two held-out instances from each class and plotted a visual comparison with the 1-Nearest Neighbor Imputation (1-NNI) baseline in Fig. 13. Notably, while MPS and 1-NNI produce similar predictions for uniformly repetitive periodic light curves, such as RR-Lyrae and contact EB (Fig. 13a, b, respectively), they diverge for less uniform patterns, such as non variable and $\gamma$ Doradus variable (Fig. 13c, d, respectively).
Appendix G Time-series classification baselines
@@
For classification performance baselines we selected three models, each representing a key category of time-series classifier; (i) distance-based, (ii) deep learning ensemble, (iii) hybrid:
@@
  • 1-NN-DTW: Nearest neighbor using a Dynamic Time Warping (DTW) similarity measure. Considered to be the most commonly used standard classical benchmark in much of the time-series classification literature [60].
@@
  • InceptionTime [61]: State-of-the-art deep learning classifier which consists of an ensemble of five Convolutional Neural Network (CNN) models.
@@
  • HC2 [62]: HC2 (HIVE-COTE V2.0) is an ensemble of classifiers, each built on different time-series data representations including phase-independent shapelets, bag-of-words based dictionaries, among others. HC2 is considered to be state-of-the-art for classification accuracy on the UCR repository.
@@
For the Astronomy dataset, we used off-the-shelf implementations of the 1-NN-DTW, InceptionTime, and HC2 classifiers that are accessible at the Python aeon (v0.11.1) library [67]. For the InceptionTime baseline, we used the recommended hyperparameter values from the original implementation [61], which are set as defaults in aeon, and trained the model for 250 epochs on an NVIDIA L40S GPU. To facilitate a fair comparison, we allowed HC2 the same computational resources as InceptionTime by setting the build-time contract to match the maximum time taken to train InceptionTime on the same hardware. We also used default hyperparameters for HC2 provided in the aeon implementation.
Appendix H Time-series imputation baselines
@@
In the main text, we compare our imputation algorithm with several state-of-the-art methods, each representing a unique architecture of time-series imputation model: (i) classical (1-NNI), (ii) diffusion model (CSDI), (iii) recurrent neural network (BRITS), and (iv) matrix completion method (CDRec):
@@
  • 1-NNI [46]: A special case of k-NNI (k-Nearest Neighbours Imputation), where $k=1$. 1-NNI substitutes missing data points in unseen time series with those from the Euclidean nearest neighbour time series in the training set.
@@
  • CSDI [56]: CSDI (Conditional Score-based Diffusion Imputation) is a generative approach to imputation which uses a neural network-based diffusion model to generate new data points that are consistent with the underlying data distribution.
@@
  • BRITS [57]: BRITS (Bidirectional recurrent imputation for time series) uses a recurrent neural network (RNN) architecture to handle missing values in time series.
@@
  • CDRec [55]: A matrix-based technique for memory efficient centroid decomposition (an approximation of SVD) of long time series.
@@
For our benchmarking investigations, we used off-the-shelf implementations of the aforementioned algorithms which are readily available in the Python PyPOTS (v0.8) library [59]. Hyperparameters for each model were set to the default values provided with the implementation.

### v2

@@ Appendix D Imputation algorithm details
Here we provide details on the specific steps of the MPSTime imputation algorithm summarized in Sec. II.3 of the main text. To elucidate the key steps of the proposed algorithm, we consider a 6-site MPS trained on time-series instances of length $T=6$ samples from some class $C$, shown in Fig. 12. Given an unseen time-series instance of data class C that is incomplete, for example, only the values $x_{2}$, $x_{4}$, $x_{6}$ are known, the goal of imputation is to estimate the unknown values (i.e., to ‘fill in the gaps’) - in this case, $x_{1}$, $x_{3}$, $x_{5}$. Our approach is performed in two main steps: (i) the trained MPS is conditioned on known values by first transforming them into quantum states and then projecting the MPS onto these known states; (ii) using the conditioned MPS, the unknown values are then determined by sequentially computing and sampling from a series of single-site reduced density matrices.
Figure 12: A toy imputation problem involving six time-series samples. We consider three observed data points (solid pink markers) $x_{2}$, $x_{4}$, $x_{6}$, and three unobserved (missing) data points, $x_{1}$, $x_{3}$, $x_{5}$ (open orange markers). Observed values in the encoding domain $x_{t}\in[a,b]$ are mapped to discrete quantum states (pink vectors) via the encoding $\phi_{i}(x_{i})$ i.e., $s_{i}=\phi_{i}(x_{i})$. The MPS is then conditioned on the observed values by making a series of projective measurements at the sites corresponding to the observed states, $s_{2}$, $s_{4}$, $s_{6}$. The unobserved states (orange vectors) are determined from the conditional probability distribution encoded by the partially conditioned MPS, and transformed back to time-series values in the encoding domain (orange markers) by an inverse transformation, $\phi_{i}^{-1}$.
@@ D.1 Projecting the trained MPS onto known states
The first step of the algorithm is to project the MPS – through a sequence of iterative site-wise measurements – into a subspace which can then be used to impute missing data. We assume the MPS $W$, which is trained to approximate the underlying joint PDF of class C time-series data, is L2 normalized. This normalization condition can be shown diagrammatically in Penrose graphical notation [52]) as:
@@
Here, we show the resulting 4-site MPS which approximates the PDF conditional on having measured the states $s_{2},s_{4}$.
@@ D.2 Imputing missing time-series data
The second step of the MPSTime imputation algorithm involves determining the values of the unobserved data points by estimating their respective values from the conditional PDF represented by the MPS. Here we use a sequential sampling approach which involves computing $d\times d$ single-site reduced density matrices.
@@
which ensures the pdf remains normalized over the support of the encoding domain $[a,b]$ ($[-1,1]$ for the Legendre basis). For details about our choice of the median as the best single point estimate, see Appendix. B.
@@
To transform the imputed states (orange vectors) back to their rescaled time-series values, one can apply the inverse of the feature map $\phi_{t}^{-1}(x)$. An additional transformation can be applied to convert the rescaled time-series (in the encoding domain) back to their respective amplitudes in the original data domain.
@@ Appendix E Inverse transform sampling
To generate time-series trajectories, as in Fig. 7, we used an inverse transform sampling (ITS) approach augmented with a rejection scheme to mitigate the effects of the encoding error discussed in Sec. B. Specifically, due to the finite basis approximation of (original continuous) time-series values, we observe an error which manifests as an artificial broadening of the conditional distribution of time-series values at each MPS site. As a consequence of this broadening, the conditional distribution can assign non-negligible probabilities to time-series values that would otherwise be highly improbable, as depicted in Fig. 10. When generating trajectories through conditional site-wise sampling – starting at the first site and proceeding sequentially to the end – these errors can propagate and compound, leading to trajectories that are not representative of the underlying joint distribution. While increasing $d$, the number of basis functions in the encoding (as in Eq. (3)), can mitigate this error in principle, achieving meaningful reductions in distribution broadening would require values that render ITS computationally intractable for most applications. Instead, we use a rejection scheme to discard highly improbable values based on their deviation from the median (i.e., weighted median absolute deviation in Eq. (31)) of the conditional distribution. Our choice of the median was motivated empirically from the analysis presented in Appendix B.
@@
Using the inverse cumulative distribution function, $F_{i}^{-1}(u)$, we select the value $x_{i}$ such that $F_{i}(x_{i})=u$. We then apply the WMAD-based rejection criteria described in Algorithm 1.
The selection of the rejection threshold factor $\alpha$ requires careful consideration. A threshold that is overly restrictive (i.e., small $\alpha$) would constrain the sampling space, potentially introducing bias by suppressing variability that is a feature of the true encoded distribution. Conversely, a less-stringent threshold (i.e., large $\alpha$) would fail to adequately address the artificial broadening effects. Through empirical investigation, we identified $\alpha=2.0$ (i.e., $2\cdot\rm WMAD)$ as an effective heuristic, which provides a balance between the two aforementioned extremes.
Algorithm 1 Single trajectory time-series generation.
@@
Appendix F Real-world time-series datasets
Here we provide further details on the three University of California, Riverside (UCR) Time Series Archive [55] datasets used for the classification task in Sec. III.3, each representing key domain of application: (i) ECG (medicine); (ii) Power Demand (industry); and (iii) Astronomy (physics). Representative time-series examples from each dataset, prior to rescaling, are shown in Fig. 13.
Figure 13: Time-series instances for each empirical dataset in the classification task. Time-series instances from each class are plotted on the same axis (and in the original data domain) for the three datasets: (a) Astronomy (2-class subset of KeplerLightCurves); (b) ECG (ECG200); and (c) Power Demand (ItalyPowerDemand). In each panel, we plot all instances from (i) class 0 (blue time-series instances), and (ii) class 1 (gold time-series instances), for each binary classification problem.
Figure 14: MPSTime imputations on unseen Kepler light curve time series with real data gaps arising from satellite transmission interruptions. The panels depict examples of MPSTime ($\chi_{\rm max}=35$, $d=12$) imputations (red) and a 1-NNI benchmark (gold) on held-out instances with real data gaps from four classes of star system: (a) RR Lyrae stars, (b) contact eclipsing binaries, (c) non-variables, and (d) $\gamma$ Doradus stars. The black points correspond to observed time-series values, while the gray horizontal lines correspond to segments of missing data, which were imputed in the original dataset [58] using the mean of the observed values.
@@ F.1 Imputation on real-world data gaps
For our investigation of the Astronomy dataset in Sec. III.2 of the main text, we excluded time-series windows containing real missing values as there exists no ground-truth for evaluation. Nevertheless, such examples provide a compelling demonstration of the real-world importance of imputation algorithms for cases when there are durations over which sensors are not recording data. For completeness, here we showcase selected examples of MPSTime imputations on real data gaps from four classes of star system (comprising the original KeplerLightCurves dataset): (i) RR Lyrae stars; (ii) contact eclipsing binaries; (iii) non-variables; and (iv) $\gamma$ Doradus stars. We trained MPSTime on length $T=100$ sample non-overlapping windows of a single time-series instance, as in Sec. III.2.2. Each example, shown in Fig. 14, compares the MPSTime imputation (red) with a 1-Nearest Neighbor Imputation (1-NNI) (gold) benchmark. Notably, while MPSTime and 1-NNI produce similar predictions for uniformly repetitive periodic light curves, such as RR Lyrae stars and contact eclipsing binaries (Figs 14a and b, respectively), the predictions of each algorithm diverge for less uniform patterns, such as non variable and $\gamma$ Doradus stars (Figs 14c and d, respectively).
Appendix G Time-series classification benchmarks
@@
For classification performance we selected three benchmarks, each representing a key category of time-series classification algorithm: (i) distance-based (1-NN-DTW); (ii) deep learning (InceptionTime); and (iii) a hybrid ensemble (HIVE-COTE v2.0). Specifically:
@@
  • 1-NN-DTW: Nearest neighbor using a Dynamic Time Warping (DTW) similarity measure. This is the most commonly used standard classical benchmark in much of the time-series classification literature [66].
@@
  • InceptionTime [67]: A state-of-the-art deep learning classifier consisting of an ensemble of five Convolutional Neural Network (CNN) models.
@@
  • HC2 [68]: HC2 (HIVE-COTE V2.0) is an ensemble of classifiers, each built on different time-series data representations, including phase-independent shapelets, bag-of-words based dictionaries, among others. At the time of the analysis, HC2 is state-of-the-art for classification accuracy on the UCR time-series repository [69].
@@
For the Astronomy dataset, we used off-the-shelf implementations of the 1-NN-DTW, InceptionTime, and HC2 classifiers that are accessible at the Python aeon (v0.11.1) library [82]. Time-series instances were $z$-score normalized (each instance along its time axis), as recommended for univariate time-series classification [55]. For the InceptionTime benchmarks, we used the hyperparameter values from the original implementation [67], which are set as defaults in aeon, and trained the model for 1500 epochs on an NVIDIA L40S GPU. To facilitate a fair comparison, we allowed HC2 the same computational resources as InceptionTime by limiting the maximum training time to match the time taken to train InceptionTime on the same hardware. We also used the default hyperparameters for HC2 as provided in the aeon implementation.
Appendix H Time-series classification training times
To compare training-time performance, we recorded the wall-clock time required to train each algorithm on the ECG dataset using representative hyperparameters from the classification benchmark in Sec. III.3. Here, we refer to “training time” as the time required to execute each classifier’s fit call on the training data. HIVE-COTE V2.0 (HC2) was excluded from the analysis since due to its algorithmic formulation, computation time results were not fairly comparable. To ensure our measurements only reflected steady-state performance, we perform one initial run for each algorithm and discard it as a burn-in. Table 4 reports the mean and standard deviation of training times over $30$ train–test resample folds. On this dataset MPSTime required on average $1.90\times 10^{1}$ s to train – nearly two orders of magnitude faster than InceptionTime ($1.72\times 10^{3}$ s), despite the latter leveraging GPU acceleration on specialized hardware (NVIDIA L40S GPU). Had InceptionTime been constrained to the same computational resources as MPSTime (i.e., a single CPU core), the training time would increase substantially. Although 1-NN-DTW remains the fastest to train ($9.90\times 10^{-3}$ s), as shown in Fig. 6 of the main text, it delivered markedly lower classification accuracy on the ECG dataset relative to all other algorithms. Taken together, our preliminary results suggest a promising speed–accuracy trade-off for MPSTime.
Table 4: Training times (mean ± std. dev.) on the ECG dataset. Training time (in seconds) across 30 resample folds of the ECG dataset is shown for the classifiers considered in our benchmark. Note that computation time results are not fairly comparable for HIVE-COTE V2.0 (HC2) due to its algorithmic formulation, and so results are not reported.
Classifier Mean (s) Std. Dev. (s) Device
[…]
HC2
§ All CPU-based timings were performed on a single core of an INTEL(R) XEON(R) GOLD 6548N CPU.
GPU timings used an entire NVIDIA L40S GPU.
Appendix I Time-series imputation benchmarks
@@
In the main text, we compare our imputation algorithm with several state-of-the-art methods, each representing different approaches to time-series imputation: (i) classical (1-NNI); (ii) diffusion model (CSDI) [62]; (iii) recurrent neural network (BRITS-I) [63]; and (iv) matrix completion method (CDRec) [61]. Specifically:
@@
  • 1-NNI [49]: A special case of $k$-NNI (k-Nearest Neighbors Imputation), where $k=1$. 1-NNI substitutes missing data points in unseen time series with those from the Euclidean nearest neighbor time series in the training set.
@@
  • CSDI [62]: CSDI (Conditional Score-based Diffusion Imputation) is a generative approach to imputation which uses a neural network-based diffusion model to infer new data points that are consistent with the underlying data distribution.
@@
  • BRITS-I [63]: BRITS-I (Bidirectional recurrent imputation for time series) is an adaptation of the BRITS algorithm to handle univariate time-series imputation. BRITS-I uses a recurrent neural network (RNN) architecture to learn a recurrent dynamical system from the observed data, and uses the learned recurrent dynamics to impute missing values in time series.
@@
  • CDRec [61] is a matrix-based technique for memory efficient centroid decomposition (an approximation of SVD) of long time series.
@@
For our benchmarking investigations in Sec. III.2.2, we used off-the-shelf implementations of the aforementioned algorithms which are readily available in the Python PyPOTS (v0.8) package [65]. Hyperparameters for each model were set to the default values provided with the implementation.